# Local Asymptotic Normality for Mixed Fractional Brownian Motion under High-Frequency Observation

### v2

@@ 6 Simulation Study
Projected (full-rank) limit.
@@
$J_{0}=0.2819517,\qquad J_{\perp}=34.17716,\qquad\mathrm{Var}(\Xi_{\sigma})=\frac{\sigma^{2}}{\pi}J_{0}=0.0897480,\qquad\mathrm{Var}(\Xi_{H}^{\perp})=\frac{\sigma^{4}}{4\pi}J_{\perp}=2.7197322.$
@@
Why projection is necessary (rank-$1$ degeneracy without it).
@@ Appendix A Closed-form expressions for $J_{0},J_{1},J_{2}$
$I(r)=\frac{1}{p}\,A^{(r+1)/p}\,\Gamma\!\left(\frac{r+1}{p}\right)\Gamma\!\left(2-\frac{r+1}{p}\right),$ (55)
@@
which gives (55) by the identity $B(x,y)=\Gamma(x)\Gamma(y)/\Gamma(x+y)$ and $\Gamma(2)=1$. Finally, differentiating under the integral sign (justified by dominated convergence for $r$ near 0) gives $I^{\prime}(0)=\int_{0}^{\infty}w(x)\ln x\,dx$ and $I^{\prime\prime}(0)=\int_{0}^{\infty}w(x)(\ln x)^{2}\,dx$. ∎
@@ A.1 Closed form for $J_{0}$
$J_{0}(H,\sigma)=\frac{2}{p}\,A^{1/p}\,\Gamma\!\left(\frac{1}{p}\right)\Gamma\!\left(2-\frac{1}{p}\right)=\frac{2}{p}\,(\sigma^{2}c_{H})^{1/p}\,\Gamma\!\left(\frac{1}{p}\right)\Gamma\!\left(2-\frac{1}{p}\right).$ (56)
@@ A.2 Closed form for $J_{1}$ and $J_{2}$
$J_{1}=C_{H}J_{0}-4K_{1},\qquad J_{2}=C_{H}^{2}J_{0}-8C_{H}K_{1}+8K_{2}.$ (57)
To express $K_{1},K_{2}$ explicitly, differentiate $\ln I(r)$ at $r=0$. From (55),
@@
$\frac{I^{\prime}(0)}{I(0)}$ $=\frac{1}{p}\Big(L_{A}+\psi(1/p)-\psi(2-1/p)\Big),$ (58)
$\frac{I^{\prime\prime}(0)}{I(0)}$ $=\Big(\frac{I^{\prime}(0)}{I(0)}\Big)^{2}+\frac{1}{p^{2}}\Big(\psi_{1}(1/p)+\psi_{1}(2-1/p)\Big).$ (59)
@@
$K_{1}=\frac{J_{0}}{2}\cdot\frac{1}{p}\Big(L_{A}+\psi(1/p)-\psi(2-1/p)\Big),$ (60)
@@
Combining (57) and (60) gives the closed forms:
$J_{1}(H,\sigma)=J_{0}(H,\sigma)\left[C_{H}-\frac{2}{p}\Big(L_{A}+\psi(1/p)-\psi(2-1/p)\Big)\right],$ (61)
@@
$J_{2}(H,\sigma)=J_{0}(H,\sigma)\left[\Big(C_{H}-\frac{2}{p}\big(L_{A}+\psi(1/p)-\psi(2-1/p)\big)\Big)^{2}+\frac{4}{p^{2}}\big(\psi_{1}(1/p)+\psi_{1}(2-1/p)\big)\right].$ (62)
@@ B.1 Step 1: remove the explicit linear term in the $H$-score
$S_{H,n}=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},\qquad R_{H,n}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}Z-\mathrm{tr}(D_{n})\Big).$ (63)
@@ B.2 Step 2: projection and orthogonalization
$R_{H,n}^{\perp}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}(D_{n}^{\perp})\Big)=R_{H,n}-\frac{\sigma a_{n}}{2}\,S_{\sigma,n}.$
@@ B.3 Singular covariance without the projection step
$\mathrm{tr}(C_{n}D_{n})$ $=\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(2L_{n}J_{0}+J_{1}+o(1)\big),$
$\mathrm{tr}(D_{n}^{2})$ $=\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(4L_{n}^{2}J_{0}+4L_{n}J_{1}+J_{2}+o(1)\big).$
@@
Write $Q(A):=Z^{\top}AZ-\mathrm{tr}(A)$. Then
@@
$\mathrm{Cov}\!\left(\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}},\frac{R_{H,n}}{L_{n}\sqrt{\mathcal{T}_{n}}}\right)=\frac{\sigma^{4}\Delta_{n}^{2p}}{\sigma L_{n}\,n\Delta_{n}}\cdot\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(2L_{n}J_{0}+o(L_{n})\big)\to\frac{\sigma^{3}}{\pi}J_{0}.$
@@ Appendix C The regime $1/2<H<3/4$
C.1 Trace approximations at order $n$.
@@
$\mathrm{tr}(C_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda+o(n),$ (64)
$\mathrm{tr}(C_{n}D_{n})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda+o(n),$ (65)
$\mathrm{tr}(D_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda+o(n).$ (66)
@@
The key point is that for $p\in(0,1/2)$ we have $f_{H},\dot{f}_{H}\in L^{2}([-\pi,\pi])$, hence $g_{n},h_{n}\in L^{2}$ uniformly because $a_{n}(\lambda)\geq 1$. In this $L^{2}$-setting, classical Szegő/Avram-type trace theorems for products of Toeplitz matrices yield the order-$n$ trace approximation with $o(n)$ remainder, see e.g. Avram [1] and standard Toeplitz references such as Grenander–Szegő [5] or Böttcher–Silbermann [2]. Applying these results to the sandwiched forms corresponding to $C_{n},D_{n}$ (equivalently, to the symbols $g_{n},h_{n}$) gives (64)–(66). ∎
@@
$\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\int_{-\pi}^{\pi}f_{H}(\lambda)^{2}\,d\lambda,$ (67)
$\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda$ $\longrightarrow\int_{-\pi}^{\pi}f_{H}(\lambda)\dot{f}_{H}(\lambda)\,d\lambda,$ (68)
$\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\int_{-\pi}^{\pi}\dot{f}_{H}(\lambda)^{2}\,d\lambda.$ (69)
@@
Since $a_{n}(\lambda)=1+\gamma_{n}f_{H}(\lambda)\to 1$ pointwise and $a_{n}(\lambda)\geq 1$, we have $|g_{n}|\leq|f_{H}|$ and $|h_{n}|\leq|\dot{f}_{H}|$. Because $f_{H},\dot{f}_{H}\in L^{2}([-\pi,\pi])$ for $p\in(0,1/2)$, dominated convergence yields (67)–(69). ∎
@@ C.3 Joint CLT for the transformed score vector and the information matrix
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}\Big(Z^{\top}C_{n}Z-\mathrm{tr}(C_{n})\Big),\qquad S_{H,n}=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},\quad R_{H,n}=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}Z-\mathrm{tr}(D_{n})\Big).$
@@
$M_{n,u}:=u_{1}\sigma^{-1}C_{n}+u_{2}\tfrac{1}{2}D_{n},\qquad Q_{n}(M_{n,u}):=Z^{\top}M_{n,u}Z-\mathrm{tr}(M_{n,u}).$
@@ Appendix D The regime $0<H<\tfrac{1}{2}$
rather than $\gamma_{n}=\sigma^{2}\Delta_{n}^{2H-1}\to\infty$. We start from the exact covariance of the observed increment vector $X_{n}=(X_{1,n},\dots,X_{n,n})^{\top}$:
$V_{n}(\sigma,H)=\mathrm{Var}(X_{n})=\sigma^{2}\Delta_{n}^{2H}\,T_{n}(H)+\Delta_{n}I_{n},$ (70)
@@
$V_{n}(\sigma,H)=\Delta_{n}^{2H}\,B_{n}(\sigma,H),\qquad B_{n}(\sigma,H):=\sigma^{2}T_{n}(H)+\varepsilon_{n}I_{n}.$ (71)
@@
$Z:=V_{n}(\sigma,H)^{-1/2}X_{n}=\Delta_{n}^{-H}\,B_{n}(\sigma,H)^{-1/2}X_{n}\sim\mathcal{N}(0,I_{n}).$ (72)
This is the same Gaussian vector $Z$ as in Lemma 1; we only rewrote the whitening in terms of $B_{n}$. Now it is natural to define
$\widetilde{C}_{n}:=B_{n}(\sigma,H)^{-1/2}\,T_{n}(H)\,B_{n}(\sigma,H)^{-1/2},\qquad\widetilde{D}_{n}:=B_{n}(\sigma,H)^{-1/2}\,\dot{T}_{n}(H)\,B_{n}(\sigma,H)^{-1/2}.$ (73)
@@
$C_{n}=\varepsilon_{n}\,\widetilde{C}_{n},\qquad D_{n}=\varepsilon_{n}\,\widetilde{D}_{n},$ (74)
@@
Assume $0<H<\frac{1}{2}$. With $Z$ as in (72) and $(\widetilde{C}_{n},\widetilde{D}_{n})$ as in (73), the score representations (9)–(11) can be rewritten as
$S_{\sigma,n}$ $=\sigma\Big(Z^{\top}\widetilde{C}_{n}Z-\mathrm{tr}(\widetilde{C}_{n})\Big),$ (75)
$R_{H,n}$ $=\frac{\sigma^{2}}{2}\Big(Z^{\top}\widetilde{D}_{n}Z-\mathrm{tr}(\widetilde{D}_{n})\Big),$ (76)
$S_{H,n}$ $=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},$ (77)
@@
Using (74) and $\gamma_{n}=\sigma^{2}/\varepsilon_{n}$, the identity (9) becomes
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}\Big(Z^{\top}(\varepsilon_{n}\widetilde{C}_{n})Z-\mathrm{tr}(\varepsilon_{n}\widetilde{C}_{n})\Big)=\sigma\Big(Z^{\top}\widetilde{C}_{n}Z-\mathrm{tr}(\widetilde{C}_{n})\Big),$
which is (75). The same argument applied to (11) gives (76). Finally, (77) is exactly (10). ∎

### v3

@@ 6 Simulation Study
6.1 Projected (full-rank) limit
@@
$J_{0}=0.2820,\quad J_{\perp}=34.1772,\quad\mathrm{Var}(\Xi_{\sigma})=\frac{\sigma^{2}}{\pi}J_{0}=0.0897,\quad\mathrm{Var}(\Xi_{H}^{\perp})=\frac{\sigma^{4}}{4\pi}J_{\perp}=2.7197.$
@@
6.2 Why projection is necessary (rank-$1$ degeneracy without it).
@@ Appendix A Closed-form expressions for $J_{0},J_{1},J_{2}$
$I(r)=\frac{1}{p}\,A^{(r+1)/p}\,\Gamma\!\left(\frac{r+1}{p}\right)\Gamma\!\left(2-\frac{r+1}{p}\right),$ (A.1)
@@
which gives (A.1) by the identity $B(x,y)=\Gamma(x)\Gamma(y)/\Gamma(x+y)$ and $\Gamma(2)=1$. Finally, differentiating under the integral sign (justified by dominated convergence for $r$ near 0) gives $I^{\prime}(0)=\int_{0}^{\infty}w(x)\ln x\,dx$ and $I^{\prime\prime}(0)=\int_{0}^{\infty}w(x)(\ln x)^{2}\,dx$. ∎
@@ A.1 Closed form for $J_{0}$
$J_{0}(H,\sigma)=\frac{2}{p{\sigma^{4}}}\,A^{1/p}\,\Gamma\!\left(\frac{1}{p}\right)\Gamma\!\left(2-\frac{1}{p}\right)=\frac{2}{p}\,{\sigma^{2/p-4}c^{1/p}_{H}}\,\Gamma\!\left(\frac{1}{p}\right)\Gamma\!\left(2-\frac{1}{p}\right).$ (A.2)
@@ A.2 Closed form for $J_{1}$ and $J_{2}$
$J_{1}=C_{H}J_{0}-4K_{1},\qquad J_{2}=C_{H}^{2}J_{0}-8C_{H}K_{1}+8K_{2}.$ (A.3)
To express $K_{1},K_{2}$ explicitly, differentiate $\ln I(r)$ at $r=0$. From (A.1),
@@
$\frac{I^{\prime}(0)}{I(0)}$ $=\frac{1}{p}\Big(L_{A}+\psi(1/p)-\psi(2-1/p)\Big),$ (A.4)
$\frac{I^{\prime\prime}(0)}{I(0)}$ $=\Big(\frac{I^{\prime}(0)}{I(0)}\Big)^{2}+\frac{1}{p^{2}}\Big(\psi_{1}(1/p)+\psi_{1}(2-1/p)\Big).$ (A.5)
@@
$K_{1}=\frac{J_{0}}{2}\cdot\frac{1}{p}\Big(L_{A}+\psi(1/p)-\psi(2-1/p)\Big),$ (A.6)
@@
Combining (A.3) and (A.6) gives the closed forms:
$J_{1}(H,\sigma)=J_{0}(H,\sigma)\left[C_{H}-\frac{2}{p}\Big(L_{A}+\psi(1/p)-\psi(2-1/p)\Big)\right],$ (A.7)
@@
$J_{2}(H,\sigma)=J_{0}(H,\sigma)\left[\Big(C_{H}-\frac{2}{p}\big(L_{A}+\psi(1/p)-\psi(2-1/p)\big)\Big)^{2}+\frac{4}{p^{2}}\big(\psi_{1}(1/p)+\psi_{1}(2-1/p)\big)\right].$ (A.8)
@@ B.1 Step 1: remove the explicit linear term in the $H$-score
$S_{H,n}=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},\qquad R_{H,n}:=\frac{\gamma_{n}}{2}\Big(Z_{n}^{\top}D_{n}Z_{n}-\mathrm{tr}(D_{n})\Big).$ (B.1)
@@ B.2 Step 2: projection and orthogonalization
$R_{H,n}^{\perp}:=\frac{\gamma_{n}}{2}\Big(Z_{n}^{\top}D_{n}^{\perp}Z_{n}-\mathrm{tr}(D_{n}^{\perp})\Big)=R_{H,n}-\frac{\sigma a_{n}}{2}\,S_{\sigma,n}.$
@@ B.3 Singular covariance without the projection step
$\mathrm{tr}(C_{n}D_{n})$ $=\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(2L_{n}J_{0}+J_{1}+{o(L_{n})}\big),$
$\mathrm{tr}(D_{n}^{2})$ $=\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(4L_{n}^{2}J_{0}+4L_{n}J_{1}+J_{2}+{o(L^{2}_{n})}\big).$
@@
Write $Q(A):=Z_{n}^{\top}AZ_{n}-\mathrm{tr}(A)$. Then
@@
$\mathrm{Cov}\!\left(\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}},\frac{R_{H,n}}{L_{n}\sqrt{\mathcal{T}_{n}}}\right)=\frac{\sigma^{4}\Delta_{n}^{2p}}{\sigma L_{n}\,n\Delta_{n}}\cdot\frac{n}{2\pi}\Delta_{n}^{1-2p}\big(2L_{n}J_{0}+{J_{1}+o(1)}\big)\to\frac{\sigma^{3}}{\pi}J_{0}.$
@@ Appendix C The regime $1/2<H<3/4$
C.1 Trace approximations at order $n$
@@
$\mathrm{tr}(C_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda+o(n),$ (C.1)
$\mathrm{tr}(C_{n}D_{n})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda+o(n),$ (C.2)
$\mathrm{tr}(D_{n}^{2})$ $=\frac{n}{2\pi}\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda+o(n).$ (C.3)
@@
The key point is that for $p\in(0,1/2)$ we have $f_{H},\dot{f}_{H}\in L^{2}([-\pi,\pi])$, hence $g_{n},h_{n}\in L^{2}$ uniformly because $a_{n}(\lambda)\geq 1$. In this $L^{2}$-setting, classical Szegő/Avram-type trace theorems for products of Toeplitz matrices yield the order-$n$ trace approximation with $o(n)$ remainder, see e.g. Avram [1] and standard Toeplitz references such as Grenander–Szegő [7] or Böttcher–Silbermann [2]. Applying these results to the sandwiched forms corresponding to $C_{n},D_{n}$ (equivalently, to the symbols $g_{n},h_{n}$) gives (C.1)–(C.3). ∎
@@
$\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\int_{-\pi}^{\pi}f_{H}(\lambda)^{2}\,d\lambda,$ (C.4)
$\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda$ $\longrightarrow\int_{-\pi}^{\pi}f_{H}(\lambda)\dot{f}_{H}(\lambda)\,d\lambda,$ (C.5)
$\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda$ $\longrightarrow\int_{-\pi}^{\pi}\dot{f}_{H}(\lambda)^{2}\,d\lambda.$ (C.6)
@@
Since $a_{n}(\lambda)=1+\gamma_{n}f_{H}(\lambda)\to 1$ pointwise and $a_{n}(\lambda)\geq 1$, we have $|g_{n}|\leq|f_{H}|$ and $|h_{n}|\leq|\dot{f}_{H}|$. Because $f_{H},\dot{f}_{H}\in L^{2}([-\pi,\pi])$ for $p\in(0,1/2)$, dominated convergence yields (C.4)–(C.6). ∎
@@ C.3 Joint CLT for the transformed score vector and the information matrix
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}\Big(Z_{n}^{\top}C_{n}Z_{n}-\mathrm{tr}(C_{n})\Big),\qquad S_{H,n}=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},$
$R_{H,n}=\frac{\gamma_{n}}{2}\Big(Z_{n}^{\top}D_{n}Z_{n}-\mathrm{tr}(D_{n})\Big).$
@@
$M_{n,u}:=u_{1}\sigma^{-1}C_{n}+u_{2}\tfrac{1}{2}D_{n},\qquad Q_{n}(M_{n,u}):=Z_{n}^{\top}M_{n,u}Z_{n}-\mathrm{tr}(M_{n,u}).$
@@ Appendix D The regime $0<H<\tfrac{1}{2}$
rather than $\gamma_{n}=\sigma^{2}\Delta_{n}^{2H-1}\to\infty$. We start from the exact covariance of the observed increment vector $X_{n}=(X_{n,1},\dots,X_{n,n})^{\top}$:
$V_{n}(\sigma,H)=\mathrm{Var}(X_{n})=\sigma^{2}\Delta_{n}^{2H}\,T_{n}(H)+\Delta_{n}I_{n},$ (D.1)
@@
$V_{n}(\sigma,H)=\Delta_{n}^{2H}\,B_{n}(\sigma,H),\qquad B_{n}(\sigma,H):=\sigma^{2}T_{n}(H)+\varepsilon_{n}I_{n}.$ (D.2)
@@
$Z_{n}:=V_{n}(\sigma,H)^{-1/2}X_{n}=\Delta_{n}^{-H}\,B_{n}(\sigma,H)^{-1/2}X_{n}\sim\mathcal{N}(0,I_{n}).$ (D.3)
This is the same Gaussian vector $Z_{n}$ as in Lemma 1; we only rewrote the whitening in terms of $B_{n}$. Now it is natural to define
$\widetilde{C}_{n}:=B_{n}(\sigma,H)^{-1/2}\,T_{n}(H)\,B_{n}(\sigma,H)^{-1/2},\qquad\widetilde{D}_{n}:=B_{n}(\sigma,H)^{-1/2}\,\dot{T}_{n}(H)\,B_{n}(\sigma,H)^{-1/2}.$ (D.4)
@@
$C_{n}=\varepsilon_{n}\,\widetilde{C}_{n},\qquad D_{n}=\varepsilon_{n}\,\widetilde{D}_{n},$ (D.5)
@@
Assume $0<H<\frac{1}{2}$. With $Z_{n}$ as in (D.3) and $(\widetilde{C}_{n},\widetilde{D}_{n})$ as in (D.4), the score representations (9)–(11) can be rewritten as
$S_{\sigma,n}$ $=\sigma\Big(Z_{n}^{\top}\widetilde{C}_{n}Z_{n}-\mathrm{tr}(\widetilde{C}_{n})\Big),$ (D.6)
$R_{H,n}$ $=\frac{\sigma^{2}}{2}\Big(Z_{n}^{\top}\widetilde{D}_{n}Z_{n}-\mathrm{tr}(\widetilde{D}_{n})\Big),$ (D.7)
$S_{H,n}$ $=\sigma\ln(\Delta_{n})\,S_{\sigma,n}+R_{H,n},$ (D.8)
@@
Using (D.5) and $\gamma_{n}=\sigma^{2}/\varepsilon_{n}$, the identity (9) becomes
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}\Big(Z_{n}^{\top}(\varepsilon_{n}\widetilde{C}_{n})Z_{n}-\mathrm{tr}(\varepsilon_{n}\widetilde{C}_{n})\Big)=\sigma\Big(Z_{n}^{\top}\widetilde{C}_{n}Z_{n}-\mathrm{tr}(\widetilde{C}_{n})\Big),$
which is (D.6). The same argument applied to (11) gives (D.7). Finally, (D.8) is exactly (10). ∎